OpenStudy (unklerhaukus):

\[\large\int\limits_0^{i\pi}e^x\,\mathrm dx\]

OpenStudy (unklerhaukus):

( nice and easy )

ganeshie8 (ganeshie8):

-2 ?

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

I would have said there's no answer...

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

can we see the area or whatever that -2 represents :o

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=%5Cint_0%5E%7Bi%5Cpi%7De%5Ex+dx not sure why wolfram is hesitating to plug e^(ipi) = 1

ganeshie8 (ganeshie8):

*e^(ipi) = -1

OpenStudy (anonymous):

I think WA is considering \(i\) to be a variable, not the square root of -1.

OpenStudy (astrophysics):

nvm

OpenStudy (unklerhaukus):

\[\begin{align} \large\int\limits_0^{i\pi}e^x\,\mathrm dx &=e^x\Big|_0^{i\pi}\\ &=e^{i\pi}-e^0\\ &=\Big[\cos(\pi)+i\sin(\pi)\Big]-1\\ &=[-1+0]-1\\ &=-2 \end{align}\]

OpenStudy (astrophysics):

Yeah :p

OpenStudy (anonymous):

A definite integral \(\displaystyle\int_a^b f(x)~dx\) is evaluated over an interval \((a,b)\), with \(a<b\). Is \(0<i\pi\)?

OpenStudy (anonymous):

\(\le\)*

OpenStudy (unklerhaukus):

i don't know what diagram could represent the integral i don't know if 0 ≤ iπ

OpenStudy (anonymous):

Regarding your second doubt, it's not - the statement doesn't make sense because the complex numbers aren't ordered.

OpenStudy (anonymous):

As for the first, you could say it's the line/contour integral over the line segment connecting the origin and (0,1) in the complex plane, but even so those limits don't make sense.

OpenStudy (unklerhaukus):

maybe 0 = iπ

OpenStudy (anonymous):

Sure, if \(i\) or \(\pi\) is 0, but we both know you mean \(i=\sqrt{-1}\) and \(\pi=3.14...\), right?

OpenStudy (unklerhaukus):

i was thinking modulo something, but yeah i = √-1 , π ~ 3

OpenStudy (unklerhaukus):

\[\begin{align} \int\limits_0^{2i\pi}e^x\,\mathrm dx &=e^x\Big|_0^{2i\pi}\\ &=e^{2i\pi}-e^0\\ &=\Big[\cos(2\pi)+i\sin(2\pi)\Big]-1\\ &=[1+0]-1\\ &=0 \end{align}\] \[\begin{align} \large\int\limits_0^{3i\pi}e^x\,\mathrm dx &=e^x\Big|_0^{3i\pi}\\ &=e^{3i\pi}-e^0\\ &=\Big[\cos(3\pi)+i\sin(3\pi)\Big]-1\\ &=[-1+0]-1\\ &=-2 \end{align}\]

OpenStudy (anonymous):

Well we have the same problem. What does it mean to integrate over an interval complex numbers with nonzero imaginary parts?

OpenStudy (anonymous):

I think I see what you're getting at, the pattern's there, but the notation and computation aren't valid imo

OpenStudy (anonymous):

Then again, replacing \(i\) with \(\sqrt{-1}\) in WA gives your answer: http://www.wolframalpha.com/input/?i=Integrate%5BExp%5Bx%5D%2C%7Bx%2C0%2CSqrt%5B-1%5D*pi%7D%5D So I don't know what to think anymore :/

OpenStudy (unklerhaukus):